# Modeling the Optical Properties of Biological Structures using Symbolic Regression

### v2

@@ I Introduction
We have previously demonstrated the potential of SR to model the optical properties of transparent and absorbing dielectrics [14] and in the context of biological photonic structures, where their optical response is inherently governed by complex physical principles [15, 16]. These studies illustrated SR’s capability to retrieve closed-form models from synthetic and experimental reflectance data. However, its application to retrieving the optical properties of biological materials remains an open challenge. To our best knowledge, no prior research has explored the use of SR in this context.
@@ II Statement of the problem
Figure 1: Biological structures studied in this work. a) Image of the inner side of a shell composed of nacre. b) SEM image of a fractured cross-section of nacre, revealing its layered structure (images a,b) are adapted from [21]). c) Photo of our Jewel Beetle’s elytron sample. d) TEM image of the elytron’s multilayer structure (adapted from [22]).
@@
with $n_{o}=\sqrt{\epsilon_{o}}$ and $n_{e}=\sqrt{\epsilon_{e}}$ being the ordinary and extraordinary refractive indices, respectively. The optical axis is the unit eigenvector corresponding to the nonrepeated eigenvalue $\epsilon_{e}$. It lies on the $xy$ plane, forming an angle $\phi$ with the $x$ axis, as indicated in Fig. 2. For an isotropic medium, $\epsilon_{o}=\epsilon_{e}=\epsilon_{h}$, and the material can be well described by its refractive index $n_{h}=\sqrt{\epsilon_{h}}$.
@@ III.1 Optical Response of a Multilayer System
×$\nabla\times\mathbf{E}=i\omega\mu_{0}\mathbf{H},\qquad\nabla\times\mathbf{H}=-i\omega\hat{\epsilon}\mathbf{E},$ (2)
@@
At each interface, the continuity of the tangential components of $\mathbf{E}$ and $\mathbf{H}$ provides the boundary conditions needed to relate the modal amplitudes across adjacent layers $n-1$ and $n$. These relations are handled using the scattering matrix (S-matrix) formalism. To properly define the S-matrix, it is necessary to reorder the amplitudes according to physical propagation direction. This involves regrouping the modes into incoming and outgoing waves at each interface, based on the sign of the real part of their $z$-component of the Poynting vector. For each pair of adjacent layers, the partial matrix $\mathbf{S}_{n-1,n}$ relates the incoming $\mathbf{A}_{in}=(A_{(n-1),1},A_{(n-1),2},A_{n,3},A_{n,4})^{T}$ and outgoing $\mathbf{A}_{out}=(A_{n,1},A_{n,2},A_{(n-1),3},A_{(n-1),4})^{T}$ modal amplitudes across the interface as
@@
$\mathbf{S}_{\text{total}}=\mathbf{S}_{0,1}\star\mathbf{S}_{1,2}\star\dots\star\mathbf{S}_{N-1,N},$ (7)
@@ III.2 Symbolic Regression Approach and Implementations
[…]
@@ IV.1 Case: Nacre Multilayer with Isotropic Aragonite
The reflectance spectra that served as the target for the SR process were numerically generated over the visible wavelength range $\lambda\in[375,800]~{}\text{nm}$. These data were smoothed using a Gaussian filter with a standard deviation $\sigma=2$ to minimize computational noise, without altering the overall spectral trends.
We performed 35 realizations, each corresponding to a different initialization seed, and the computing time per realization was about 1.9 hours. This expanded dataset provided a statistically significant set of predicted models to analyze the consistency of the retrieved expressions. The best retrieved models for the refractive index $n_{h}^{SR}$ are summarized in Table 1.
@@
For the sake of completeness, we also compare the results predicted by Eq. 11 with those obtained with a Deep Neural Network (NN). For this numerical experiment, the network maps full reflectance spectra to the corresponding refractive index spectra over the visible range. It consists of an input layer with $N_{\lambda}$ neurons, three hidden layers with 256, 128, and 64 neurons, respectively, and an output layer with $N_{\lambda}$ neurons, where $N_{\lambda}$ denotes the number of sampled wavelengths. To train the NN, we first generated a synthetic database comprised of 10.000 samples, each corresponding to a pair of a possible refractive index value with its associated reflectance spectrum. The generation of the database required nearly 2 hours of computation. The network was then trained on this dataset, which took 1.14 hours. After training, the model was used to predict the refractive index $n_{h}^{\text{NN}}$. In contrast, the SR approach required only a single reflectance spectrum as training, avoiding the need for a precomputed database.
@@
In Fig. 4-b) we compare the reflectance spectra related to the SR model and the target reflectance data. As observed, the response of the SR model has a high prediction accuracy with a $\text{RMSE}_{R}=0.01109$ metric value. Despite the small differences present for wavelengths below $0.4\;\mu$m, the overall optical response predicted by the SR agrees with the target spectrum.
To further understand the operating principles of the SR approach, we repeated the previous numerical experiment keeping the same conditions. However, this time we searched for the closed-form expression of the dielectric function $\epsilon$ instead of searching for that of the refractive index $n$. This numerical experiment took about 1.3 hours per realization.
@@
The results from this first case study are encouraging, as they demonstrate the ability of SR to capture intrinsic relationships in the data. Furthermore, the underlying physical relation $\epsilon_{h}=n_{h}^{2}$ is still preserved. This highlights the robustness and consistency of our SR approach.
@@ IV.2 Case: Nacre Multilayer with Anisotropic Aragonite
We used numerically generated spectra as input data. Also, to avoid a significant increase in computing time, instead of the $N_{n}=100$ periods initially considered we fixed them to $N_{n}=10$. As in the first case study, we performed 35 realizations corresponding to different random seeds and the total computing time for each realization was about 3.9 hours.
@@ IV.3 Case: Chrysochroa Jewel Beetle Multilayer
We conducted a numerical experiment analogous to the previous cases, performing 35 realizations of the SR process. The total computation time for each realization was approximately 1.23 hours.

### v3

@@ I Introduction
We have previously explored the potential of SR to model the optical properties of transparent and absorbing dielectrics [14] and also of transparent biological materials such as keratin and chitin [15]. The results reported in those references showed the ability of SR to recover closed-form models from synthetic reflectance data. However, due to the operating principles of the numerical library used (see [14] for details), they also made evident the impossibility to take into account the dimensional homogeneity and the algebraic complexity of the expressions searched. We have recently used the PhySO numerical library [12] to model the spectral reflectance of photonic biological structures made of disordered materials [16]. The results reported in this reference settled the basis for the use of SR to model the optical properties of biological photonic structures. To the best of our knowledge, the performance of SR has not been explored in this context before. This presents a formidable challenge due to the complexity of biological structures.
@@ II Statement of the problem
Figure 1: Biological structures studied in this work. a) Image of the inner side of a shell composed of nacre. b) SEM image of a fractured cross-section of nacre, revealing its layered structure (images (a),(b) are adapted from Ref. [21]). c) Photo of our Jewel Beetle’s elytron sample. d) TEM image of the elytron’s multilayer structure (adapted from Ref. [22]).
@@
with $n_{o}=\sqrt{\epsilon_{o}}$ and $n_{e}=\sqrt{\epsilon_{e}}$ being the ordinary and extraordinary refractive indices, respectively. The optical axis is the unit eigenvector corresponding to the nonrepeated eigenvalue $\epsilon_{e}$. For the purpose of keeping the complexity of the problem at a manageable level, we will assume that the optical axis is located in the $xy$ plane, forming an angle $\phi$ with the $x$-axis, as indicated in Fig. 2. However, it is important to note that this simplified configuration was chosen for illustration purposes only and does not correspond to the general case of anisotropic biological materials that can be found in nature. For an isotropic medium, $\epsilon_{o}=\epsilon_{e}=\epsilon_{h}$, and the material can be well described by its refractive index $n_{h}=\sqrt{\epsilon_{h}}$.
@@ III.1 Optical Response of a Multilayer System
$\nabla\times\mathbf{E}=i\omega\mu_{0}\mathbf{H},\qquad\nabla\times\mathbf{H}=-i\omega\hat{\epsilon}\mathbf{E},$ (2)
@@
At each interface, the continuity of the tangential components of $\mathbf{E}$ and $\mathbf{H}$ provides the boundary conditions needed to relate the modal amplitudes across adjacent layers n1n-1 and $n$. These relations are handled using the scattering matrix (S-matrix) formalism. To properly define the S-matrix, it is necessary to reorder the amplitudes according to physical propagation direction. This involves regrouping the modes into incoming and outgoing waves at each interface, based on the sign of the real part of their $z$-component of the Poynting vector. For each pair of adjacent layers, the partial matrix $\mathbf{S}_{(n-1),n}$ relates the incoming $\mathbf{A}_{in}=(A_{(n-1),1},A_{(n-1),2},A_{n,3},A_{n,4})^{T}$ and outgoing $\mathbf{A}_{out}=(A_{n,1},A_{n,2},A_{(n-1),3},A_{(n-1),4})^{T}$ modal amplitudes across the interface as
@@
$\mathbf{S}_{\text{total}}=\mathbf{S}_{0,1}\star\mathbf{S}_{1,2}\star\dots\star\mathbf{S}_{(N-1),N},$ (7)
@@ III.2 Symbolic Regression Approach and Implementations
Specify Allowed Operations and Constants: The next step involves defining the building blocks that PhySO is allowed to use for generating the predicted models $n^{SR}$. This is a set of mathematical operations, which could include arithmetic operations (e.g., addition, subtraction, multiplication, division), exponentials, logarithms, or trigonometric functions. Additionally, PhySO requires constants with specified dimensions to be included in the search space. The choice of the allowed building blocks is entirely dependent on the problem and the nature of the search. Since the materials considered in this work are non-absorbing dielectrics, the four basic operations ${+,,,/}\{$+,-,*,/\} are sufficient to achieve our objectives. This choice was confirmed by extensive numerical experiments, not included here due to space limitations. These experiments showed that the SR could discriminate unnecessary operations and tended to retain only the basic ones in the final expressions. Regarding the allowed constants, our rule of thumb is to use two constants for each variable with a matching dimension. In this work, we defined the set of allowed constants as $\{A(\mu\text{m}),B(\mu\text{m}),C,D\}$, where the first two match the units of $\lambda$, and the other two are dimensionless as the response. Note that there is an optimization process to find the numerical value of each one, therefore, each allowed constant adds considerable computing time.
@@ IV.1 Case: Nacre Multilayer with Isotropic Aragonite
The reflectance spectra that served as the target for the SR process were numerically generated over the visible wavelength range $\lambda\in[0.375,0.800]~\mu\text{m}$. These data were smoothed using a Gaussian filter with a standard deviation $\sigma=2$ to minimize computational noise, without altering the overall spectral trends.
We performed 35 realizations, each corresponding to a different initialization seed, and the computing time per realization was about 1.9h. This expanded dataset provided a statistically significant set of predicted models to analyze the consistency of the retrieved expressions. The best retrieved models for the refractive index $n_{h}^{SR}$ are summarized in Table 1.
@@
For the sake of completeness, we also compare the results predicted by Eq. 11 with those obtained with a Deep Neural Network (NN). For this numerical experiment, the network maps full reflectance spectra to the corresponding refractive index spectra over the visible range. It consists of an input layer with $N_{\lambda}$ neurons, three hidden layers with 256, 128, and 64 neurons, respectively, and an output layer with $N_{\lambda}$ neurons, where $N_{\lambda}$ denotes the number of sampled wavelengths. To train the NN, we first generated a synthetic database comprised of 10000 samples, each corresponding to a pair of a possible refractive index value with its associated reflectance spectrum. The generation of the database required nearly 2h of computation. The network was then trained on this dataset, which took 1.14h. After training, the model was used to predict the refractive index $n_{h}^{\text{NN}}$. In contrast, the SR approach required only a single reflectance spectrum as training, avoiding the need for a precomputed database.
@@
[…]
To further understand the operating principles of the SR approach, we repeated the previous numerical experiment keeping the same conditions. However, this time we searched for the closed-form expression of the dielectric function $\epsilon$ instead of searching for that of the refractive index $n$. This numerical experiment took about 1.3h per realization.
@@
The results from this first case study are encouraging, as they demonstrate the ability of SR to capture intrinsic relationships in the data. Furthermore, the underlying physical relation $\epsilon_{h}=n_{h}^{2}$ is still preserved. This highlights the robustness and consistency of our SR approach. However, we also found that expressions retrieved by SR with a physical meaning do not necessarily represent the optimal solution that minimizes the chosen regression metric. This interesting finding will repeatedly appear throughout the numerical experiments conducted in this study.
@@ IV.2 Case: Nacre Multilayer with Anisotropic Aragonite
We used numerically generated spectra as input data. Also, to avoid a significant increase in computing time, instead of the $N_{n}=100$ periods initially considered we fixed them to $N_{n}=10$. As in the first case study, we performed 35 realizations corresponding to different random seeds and the total computing time for each realization was about 3.9h.
@@ IV.3 Case: Chrysochroa Jewel Beetle Multilayer
We conducted a numerical experiment analogous to the previous cases, performing 35 realizations of the SR process. The total computation time for each realization was approximately 1.23h.
@@
Despite its simplicity, this last case study clearly illustrates the usefulness of SR when only experimental measurements are available. The results demonstrate that it is possible to obtain a closed-form expression that predicts the property under study within the considered spectral range, even with limited input data. It should be noted that this example is more complex because the model of the property being studied is obtained indirectly from the measured response.
On the other hand, this case study also reveals two important flaws of SR that should be considered when using it. First, while the final expression will be readable and dimensionally homogeneous, it may not have physical meaning because this characteristic does not imply the best fit of the chosen metric. Second, SR depends entirely on the geometric model used to represent the structure under study and the method used to calculate the response that fits the experimental data. In the presented case study, the multi-layered structure with flat interfaces is clearly a simplified representation of the beetle wing structure under consideration. Therefore, this lack of information inevitably influences the algebraic form of the final expression.